Astronomy Report – December 5th

Name:  Thomas Horn

Crew:  184

Date:  12/5/2017

Sky Conditions:  Clear

Wind Conditions:  Still

Observation Start Time:  8:30am

Observation End Time:  9:30am

Summary:  Did a scan of solar features for unusual activity.  None
observed.  Image of solar surface attached.

Objects Viewed:  Sun

Problems Encountered:  None

Sol Summary – December 5th

Crew 184 Sol Summary Report  05 Dec 2017

MDRS Sol Summary Report for Sol 4

Summary Title:  Our first failure and major replan

Mission Status:  Today was a hectic day with three major failures our
team dealt with.
1.    Internet – Between 8:30am and 9:30am 800MB of bandwidth was
utilized. It is unknown what caused this but likely culprit was a
device being left on and doing a software update in the background.
Internet communication was lost at this time.  These reports are being
delivered by Shannon today.
2.    Water – During our morning fill of the loft tank the pump
was non-functional.  Possible causes were quickly narrowed down to
either a frozen pipe due to the below freezing temperatures or a
clogged filter. Our team immediately switched to using bottled water
which we keep as an emergency reserve while we troubleshot the issue.
At 1pm we called Mission Control for advice who recommended we connect
the pipe heating system and wait till midday, this resolved the issue
around 3pm and water was restored to the HAB
3.    Power – During our morning check of SOC it was noted that
the generator was not running, having failed sometime during the
night, and that SOC was at 3%.  We immediately instituted a shutdown
of MDRS powered equipment to preserve what little power we had and
investigated the generator.  It was noted that a significant oil leak
had developed from the oil cap, upon tightening the cap the leak
stopped and the generator was restarted.  When returning to the
science dome it was observed that SOC was hovering around 10% and did
not appear to be charging.  After consultation with Mission Control it
was discovered that the generator was in idle due to the power system
not requesting it to go to full delivered power given the fact that
the solar array system was charging the battiers (however slowly) .
In addition the SOC was likely incorrect due to it being a calculated
value and us confusing it due to going off scale.  We will assess SOC
tonight after a full day in the sun and decide forward plan.

Our morning was mostly consumed with maintenance and troubleshooting,
we quickly discovered that when receiving failures our level of
knowledge of the systems was exhausted in how to troubleshoot,
recommend future crew get access to and study manuals for MDRS
equipment prior to stay.

In the afternoon when we got our facilities in a stable configuratrion
four members of our team went on our first Matryoshka EVA. We were
able to go to the Kissing Camel ridge and collect two samples, however
we were unable to go to the other site in URC South due to delays in
getting out since we had to fix our non-nominal systems.

The two sample we collected were to highlight interesting
geomorphological features to be used in outreach activities back in
Oxford. The strike and dip of the ridge component the first sample was
extracted from was also taken. The intention of sample collection and
strike and dip measurement is to determine whether satellite and
outcrop scale imagery data is reliable at predicting these.

Sol Activity Summary:
1.    Troubleshooting and EVA as discussed above
2.    Exercise
3.    Assembling of exercise equipment using martian dirt
(medicine ball complete!)
4.    HAB Cleaning
5.    Solar observation

Look Ahead Plan:
With internet down we are unable to contact our scheduling team.  We
will receive the schedule in the morning when internet bandwidth
restarts and go over it as a team to plan our day.

Anomalies in work:
1.    Yamaha 250 ATV #3 wheel deemed non-recoverable.  We have
parked it off to the side for mission support.  Photos of damage

Weather: Very cold!  It was below freezing for most of the day.

Crew Physical Status: All in good health.

EVA: Delayed due to non-nominal systems, so we are requesting the same
EVA from Sol 3 but just for site A (mound on URC South) to finish off
the sample collection. The plan and crew utilised will be the exact
same as today.

Reports to be file:
EVA Request
Sol Summary

Support Requested:
Trash placed in the Rear Airlock for pickup.  Burnable trash on the
side with the box.  Non-burnable trash on the other side.

Operations Report – December 5th

Crew 184 Operations Report 12/05/2017


Name of person filing report: Akash Trivedi

Non-nominal systems: Water (pump/lines – static to loft), generator
failure, sudden loss of data

Notes on non-nominal systems:  Water: tanks from static to loft not
transferring due to speculated frozen pipe or blockage. Waited until
10am as per protocol to see if frozen pipes thawed. At 10am, still
frozen. Waited until 12pm for sunlight to be in contact with pipes to
see if thaws. After the meeting at lunch, we learned that a plug by
the static tank needed to be seated. This allows heating along the
pipes. Once plugged we waited some time, and pump was operational
again at around 3pm. Anomaly resolved.

Generator: Noticed generator off at 0900 having failed during the
night. Turned on and seemed to be operating nominally. Returned to
check up on it at 1000, and noticed a coolant leak from the radiator.
Generator turned off and leak corrected by tightening the antifreeze
cap. Generator turned back on and checked status on power console –
seems to not be doing the charging as it’s on “absorb” and SOC not
increasing despite minimal hab power usage. After Lunch meeting, we
monitored thegenerator closely and at sunset the SOC returned to 100%.
Will continue to monitor closely, but power is good at this time.

Loss of internet data: Data check was conducted at 0845 when it was
approximately 800MB, and an hour later, this dropped to zero. Reason

Generator (hours run):  Unknown due to night time failure

Solar— SOC % (Before generator is run at night) 4% at 0900, 14% at
1000 (when coolant leak discovered), static 9% on “absorb charging”
with minimal hab usage and generator powered on

Diesel – 48%

Propane –  79 percent volume

Ethanol Free Gasoline (5 Gallon containers for ATV) – 1.25 gallons

Water (trailer) – 600 gallons

Water (static) – 450 gallons

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 46 gallons

Toilet tank emptied: Yes

ATV’s Used: Honda ATV, Yamaha 1

Oil Added? No

ATV Fuel Used: 1.25 gallons

# Hours the ATVs were Used today: 00:45

Notes on ATVs: Yamaha 3 Rear right tire is still wobbly, EVA team
assessed and determined it will require. Photos will be provided in
overall Sol 4 package for awareness and review.

Deimos rover used: Used

Hours: 00:45

Beginning charge: 100%

Ending charge: 100%

Currently charging:Yes

Sojourner rover used:  ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used:  Not used


Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used:  Not Used


Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:  Not Used


Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where?  No

General notes and comments: Multiple non-nominal systems!!!

Summary of internet: Sudden anomalous data loss – see above

Summary of suits and radios: All nominal based on last usage

Summary of Hab operations: Loft water tank almost empty, multiple
non-nominal systems

Summary of GreenHab operations:  All nominal

Summary of ScienceDome operations: All nominal

Summary of RAMM operations: Not Operational

Summary of health and safety issues: No problems now that water, heat,
and power are back.

Questions, concerns and requests to Mission Support: See above on
non-nominal systems. Lack of training and relevant manuals for
troubleshooting significant issues in mission critical systems. Poor
insulation of static tank to loft tank pipes – could lead to pipe

GreenHab Report – December 5th

GreenHab Report

Crewmember’s name: Akash Trivedi

Date: 5/12/17

Environmental control:

Ambient (no heating or/cooling): No wind, door open at 9am. Since
generator failed and heater would not work, we closed door to
stabalise temperature.

Heating: Nominal, once generator was back on

Cooling: Not operational

Both heating and cooling

Shade cloth on

Average temperature: 21 oC in morning before power shortage, 16 oC at 1700

Changes to crops:  Same tomato plant droopy, also noticed a droopy
radish plant and moved it to a brighter location

Daily water usage for crops: Nominal

Time(s) of watering for crops: 9am

Research observations: (Currently not operational)

Changes to research plants: N/A

Daily watering and amount of water used: 2 gallons

Aquaponics: (Currently not operational)

Narrative: Plants are doing good on the whole!

Support/supplies needed: Fire extinguisher required, extra mineral
salts for supplements in case of deficiencies

Journalist Report – December 5th

Crew 184 Journalist Report

Willie Schumann

05 December 2017

Title                            Crisis Management

Narrative                  Some people say you only really arrived to a new home, when you managed your first crisis and you didn’t run away. Well, that means we have finally arrived on Mars. We woke up to a bunch of problems, which could have had a dramatic effect on the livelihood of crew number 184.

Crew Engineer Hunt woke up early today and was the first to realize, that our internal water tank in our top floor had reached a critical low level. This tank fuels our kitchen, the bath and the toilet – it is the heart of our Martian shelter. It is just below the roof, because it uses gravity to deliver water to the various outlets below. At first we thought, our pump was defective, that enables the whole operation.

That conclusion was plausible, because shortly after crisis number one we detected crisis number two. Our battery, which is fueled by a generator and the solar panels at their respective working hours, was down to five percent. Commander Horn and Crew Engineer Hunt, still in his pajamas, observed the devices and detected a leaking oil tank.

Was our failing energy source the reason for the block in the water delivery? And could the Officer Hunt fix the oil leak and subsequently fix the hab? Time was crucial, because we only had limited reserve water bottles and the toilet was also relying on the tank. From personal experience I could tell, holding back on business at the space-loo makes it impossible to stay operational.

The night was awfully cold leading to Sol 4 and so we came up to a new possible reason of the failing water system. Maybe the pipes leading into the hab were frozen? We checked them and although we couldn’t peek inside we could feel how cold they were. The sun still hadn’t turned around far enough to warm up the external water tank. We never had been so anxious for the hot giant star to move faster to hit Mars with a wave of heat.

After two hours of trouble shooting and learning more about our life-sustaining infrastructure we contacted mission control and requested assistance on our problems. We were assured, that we were on the right track and that we will life another day. After we closed the oil leak, which was caused by cap, which wasn’t screwed tight enough, the battery gained quickly power. And once the sun turned the water flooded into our hab.

Because of the crisis situation we had to push our second EVA’s two to three hours back. Me and a crew of three, led by Science Officer Akash Trivedi, were already in the pre-breathing chamber, when we heard the good news of the resolved problems. We could start with a light heart to our mission to collect soil and rock samples.

Akash Trivedi is one of to European members of our crew. The Briton is well connected to the university of Oxford, which asked him to do a so called Matryoshka project. He received satellite data for interesting surfaces on Mars and now wants to collect samples from exactly these spots. Like a Russian doll both elements will complete one another.

For the first time we took the rover out and as we reached our destination climbed on hills to collect the sources. It was fun. The heavy helmet and backpack didn’t really hold me back. But I must say carrying the camera equipment and the necessity to be faster at certain spots and staying longer to have enough flesh for the footage is demanding. It sounds contradicting, but creating great movie footage makes often only sense for the filmmaker.

Coming back was great, because there was a special treat waiting for us. With restored energy and water, we will have our first shower on Mars. We were holding back on it since the beginning, looking on the water crisis, quite a good exercise to deal with such a shortage. So if you excuse me, I have a date with our shower…

Personal Logbook             Today was a fruitful day, but I will be happy when I lay down to sleep.  Partly this is due to our busy schedule, to the high demand of the marsonauts life and the challenges of the work.

Because my crew has little to no experience with documentary work, it is also a workshop for them to align with the demands of filmmaking. In a way, they have to be actors, but actors, who play themselves. Once you understand this and the technical requirements of filmmaking you are half way there.

Some are very talented and pure gold. For others I have to explain certain procedures over and over again. And it’s all good and no ones fault. It just makes it a bit harder for me. Which is tough, when you already do four jobs by yourself.

Today I could also do some more photos on film, which made me very content. During the various problems we had to manage I was doing some photography of our polished helmets and suits. I thought I use the calm to cover theses essentials.

Tomorrow I might stay in the hab for a day. I definitely should take a little rest in one of the next days. I think I will make this decision upon the weather forecast. Still super excited to be here, but taking care of my energy level to continue to deliver good pictures.

Thank you very much for your help and attention.

Willie Schumann, Journalist, Crew 184

Copyright © The Mars Society. All rights reserved. | Main Site